# Do physics-informed neural networks (PINNs) need to be deep? Shallow PINNs using the Levenberg–Marquardt algorithm

### v1

@@ 3.4 Inverse problem: three-dimensional Bratu equation
For training, we randomly select $n=2000$ collocation points within the domain. The corresponding reference solution is obtained using the symmetric finite difference method (SFDM) [35] with 300 uniform subintervals (301 grid points) in each spatial direction. The true parameters are set to $\lambda_{1}=C=2$ and $\lambda_{2}=1$. The identification is performed using the architecture NN(3,25,25,1)(3,25,25,1), resulting in a total of 778 trainable parameters, including the network weights, biases, and the unknown parameters $\lambda_{1}$ and $\lambda_{2}$.
@@
Fig. 5(d) presents the training loss curves for both optimization algorithms. LM achieves a final loss value of $5.7\times 10^{-10}$, which is substantially lower than the $3.8\times 10^{-5}$ obtained by BFGS. In terms of computational cost, LM requires approximately 2470 seconds, which is slightly higher than the 1851 seconds required by BFGS.
@@
LM $\Delta u+1.99994e^{1.00018u}$
BFGS $\Delta u+2.01510e^{0.97944u}$
@@
LM $5.7\cdot 10^{-10}$ 0.0028 0.0180 2470
BFGS $3.8\cdot 10^{-5}$ 0.7549 2.0562 1851
@@
NN(3,5,5,1)(3,5,5,1) $1.4\cdot 10^{-4}$ 2.0648 77.2285 87
NN(3,10,10,1)(3,10,10,1) $1.2\cdot 10^{-7}$ 0.0169 0.7576 211
NN(3,15,15,1)(3,15,15,1) $1.4\cdot 10^{-7}$ 0.0213 0.0337 404
NN(3,20,20,1)(3,20,20,1) $2.4\cdot 10^{-8}$ 0.0090 0.1370 1051
NN(3,25,25,1)(3,25,25,1) $5.7\cdot 10^{-10}$ 0.0028 0.0180 2470
Finally, Tab. 9 compares the loss values, APEs, and computational times for various network architectures trained using LM. Increasing the number of neurons generally improves both accuracy and loss values. Notably, even the smaller network NN(3,10,10,1)(3,10,10,1) yields reasonably accurate results, with APEs for all parameters remaining below 1%, and also outperforms BFGS (see Tab. 8) in terms of both accuracy and computational time.
@@ 4 Conclusion
Numerical experiments are conducted for both forward and inverse problems involving the Burgers, Schrödinger, Allen–Cahn, and three-dimensional Bratu equations. Using a fixed architecture NN(a,25,25,1)(a,25,25,1) in all cases, LM consistently outperforms BFGS. Specifically, LM achieves final loss values on the order of $10^{-8}$ to $10^{-10}$, corresponding to improvements of three to five orders of magnitude over BFGS. Moreover, LM reaches this accuracy with substantially fewer iterations (4000 versus 20000). In forward problems, LM also produces relative $L_{2}$ errors that are approximately two orders of magnitude smaller, indicating a more accurate approximation of the PDE solutions. In inverse problems, LM not only yields lower loss values but also converges more rapidly and accurately to the true parameter values, highlighting its effectiveness for parameter identification.
@@
Overall, our results indicate that a wide class of PDEs can be efficiently solved using shallow PINNs when the problem is reformulated as a nonlinear system and solved with LM. For practical applications, we recommend starting with shallow architectures such as NN(a,10,10,1)(a,10,10,1) or NN(a,15,15,1)(a,15,15,1), which offer a favorable balance between accuracy and computational cost, and increasing the number of neurons only when the problem exhibits complex features such as sharp gradients or multiscale behavior.

### v2

@@ 3.4 Inverse problem: three-dimensional Bratu equation
For training, we randomly select $n=10000$ collocation points within the domain. The corresponding reference solution is obtained using the symmetric finite difference method (SFDM) [35] with 300 uniform subintervals (301 grid points) in each spatial direction. The true parameters are set to $\lambda_{1}=C=2$ and $\lambda_{2}=1$. The identification is performed using the architecture NN(3,25,25,1)(3,25,25,1), resulting in a total of 778 trainable parameters, including the network weights, biases, and the unknown parameters $\lambda_{1}$ and $\lambda_{2}$.
@@
Fig. 5(d) presents the training loss curves for both optimization algorithms. LM achieves a final loss value of $9.4\times 10^{-7}$, which is substantially lower than the $1.2\times 10^{-3}$ obtained by BFGS. In terms of computational cost, LM requires approximately 5802 seconds, whereas BFGS takes about 16262 seconds.
@@
LM $\Delta u+1.99992e^{1.00040u}$
BFGS $\Delta u+1.99144e^{1.05985u}$
@@
LM $9.4\cdot 10^{-7}$ 0.0041 0.0401 5802
BFGS $1.2\cdot 10^{-3}$ 0.4281 5.9845 16262
@@
NN(3,5,5,1)(3,5,5,1) $2.7\cdot 10^{-2}$ 5.1426 71.8921 286
NN(3,10,10,1)(3,10,10,1) $2.9\cdot 10^{-3}$ 0.9318 12.8455 831
NN(3,15,15,1)(3,15,15,1) $2.0\cdot 10^{-4}$ 0.0560 0.8668 2505
NN(3,20,20,1)(3,20,20,1) $4.9\cdot 10^{-5}$ 0.0212 0.3051 4255
NN(3,25,25,1)(3,25,25,1) $9.4\cdot 10^{-7}$ 0.0041 0.0401 5802
Finally, Tab. 9 compares the loss values, APEs, and computational times for various network architectures trained using LM. Increasing the number of neurons generally improves both accuracy and loss values. Notably, even the smaller network NN(3,15,15,1)(3,15,15,1) yields reasonably accurate results, with APEs for all parameters remaining below 1%, and also outperforms BFGS (see Tab. 8) in terms of both accuracy and computational time.
@@ 4 Conclusion
Numerical experiments are conducted for both forward and inverse problems involving the Burgers, Schrödinger, Allen–Cahn, and three-dimensional Bratu equations. Using a fixed architecture NN(a,25,25,1)(a,25,25,1) in all cases, LM consistently outperforms BFGS. Specifically, LM achieves final loss values on the order of $10^{-7}$ to $10^{-10}$, corresponding to improvements of three to five orders of magnitude over BFGS. Moreover, LM reaches this accuracy with substantially fewer iterations (4000 versus 20000). In forward problems, LM also produces relative $L_{2}$ errors that are approximately two orders of magnitude smaller, indicating a more accurate approximation of the PDE solutions. In inverse problems, LM not only yields lower loss values but also converges more rapidly and accurately to the true parameter values, highlighting its effectiveness for parameter identification.
@@
Overall, our results indicate that a wide class of PDEs can be efficiently solved using shallow PINNs when the problem is reformulated as a nonlinear system and solved with LM. For practical applications, we recommend starting with shallow architectures such as NN(a,15,15,1)(a,15,15,1), which offer a favorable balance between accuracy and computational cost, and increasing the number of neurons only when the problem exhibits complex features such as sharp gradients or multiscale behavior.